OpenStudy (anonymous):

pete is 14 and his grandfather is 54. how many years ago was his grandfather 6 times as old as pete? step by step plz

OpenStudy (anonymous):

Difference in age = 40 Let pete's previous age be x Let g.father's previous age be 6x 6x - x = 40 5x = 40 x = 8 Hence when Pete was 8 the condition was met. He was 8 years old, 6 years ago. So I think the answer is 6.

OpenStudy (anonymous):

I've got the same answer but a different method. Put (grandpa's age - x) = 6(pete's age - x) and solve for x, where x is the number of years ago that grandpa was 6 times older than pete. so, 54 - x = 6(14 - x)...

OpenStudy (anonymous):

From Mathematica:\[\text{Solve}[\{6(p-x)==(g-x)\text{/.} \{p\to 14,g\to 54\}\},x] \]\[\{\{x\to 6\}\} \]